Assume that a sample is used to estimate a population proportion μ . Find the margin of error M.E. that corresponds to a sample of size 408 with a mean of 39.1 and a standard deviation of 11.3 at a confidence level of 99.9%. Confidence Interval for Means Calculator

Report ME accurate to one decimal place because the sample statistics are presented with this accuracy. M.E. =

Answer should be obtained without any preliminary rounding. However, the critical value may be rounded to 3 decimal places.
Answer #1

Solution :

Given that,

Sample size = n = 408

Z/2 = 3.291

Margin of error = E = Z/2* ( /n)

= 3.291 * (11.3 / 408)

Margin of error = E = 1.8